# Unsupervised deep learning model for fast energy layer pre-selection of delivery-efficient proton arc therapy plan optimization of nasopharyngeal carcinoma

### v1

@@ Abstract
Objective. Proton arc therapy (PAT) is an emerging and promising modality in radiotherapy, offering several advantages over conventional intensity-modulated proton therapy (IMPT). However, identifying the optimal energy layer (EL) sequence remains computationally intensive due to the large number of possible energy layer transitions. This study proposes an unsupervised deep learning framework for fast and effective EL pre-selection, aiming to minimize energy layer switch time while preserving high plan quality. Approach. We introduce a novel data representation method—spot-count representation—which encodes the number of proton spots intersecting the target and organs at risk (OARs) in a matrix structured by sorted gantry angles and energy layers. This representation is the input of a U-Net-based architecture, SPArcdl, which is trained to optimize a tri-objective function: maximizing target coverage, minimizing OAR exposure, and reducing energy switching time. The model is evaluated on $54$ nasopharyngeal cancer cases, and its performance is benchmarked against plans generated by SPArcparticle swarm. Main results. SPArcdl produces EL pre-selection that significantly improves both plan quality and delivery efficiency. Compared to SPArcparticle swarm, it enhances the conformity index by $0.16$ ($p<0.01$), reduces the homogeneity index by $0.71$ ($p<0.01$), shortens the energy switching time by $38.4\%$ ($p<0.01$), and lowers the mean dose to brainstem by $0.21$ ($p<0.01$). The results unintentionally reveal employing unchanged ELS is more time-wise efficient than descended ELS. SPArcdl’s inference time is within 1 second. Significance. SPArcdl is a fast and effective tool for generating high-quality PAT plans by strategically pre-selecting energy layers to reduce delivery time while maintaining excellent dosimetric performance. The proposed spot-count representation lays a foundation for incorporating unsupervised deep learning approaches into the optimization of proton arc therapy planning.
@@ 1 Introduction
A major approach to solve the PAT problem is to break it down into two steps: pre-selecting ELs in advance, then using the selected ELs to complete regular spot weight optimization [27]. A significant challenge to the clinical adoption of PAT lies in the optimization of energy layer (EL) sequences. In SPArc, the system is required to deliver protons at selected energy layers on each gantry angle. With hundreds of control points involved, frequent switching between energy layers can lead to substantial delays in treatment delivery. This is especially true for upward energy switches, which are time-intensive due to the magnetic hysteresis in the proton therapy system [8, 30, 2, 23, 22]. Consequently, minimizing energy layer switching (ELS) time without compromising dosimetric quality has become a focus in PAT plan optimization.
@@
Several EL sequence optimization approaches have been proposed to address this challenge. Sanchez-Parcerisa et al. [19] proposes and compares four range optimization algorithms to select optimal mono- and bi-energetic beam energies for proton modulated arc therapy (PMAT), aiming to minimize energy-switching time while maintaining dosimetric quality. However, this method shows constraints in performance when applied to intricate or irregular target geometries. The SPArc algorithm [6] improves dose delivery by iteratively refining gantry angles and redistributing ELs. As a subsequent method, SPArcseq [12] reduces switching time by aligning energy sequences with the gantry’s rotational direction, and SPArcparticle swarm (SPArcps) [17] builds upon it by segmenting the arc into sectors based on mean water equivalent thickness (WET) and selecting ELs that minimize switching time across control points. However, SPArcps may need more objective functions on organs at risk (OARs) to optimize Head and Neck cancer, including nasopharyngeal carcinoma (NPC). Additionally, the ELO-SPAT framework [28] integrates multiple objectives, including dose fidelity, energy sequencing constraints, and regularization terms, into a unified mathematical formulation that penalizes upward energy switches while exploiting downward transitions to enhance delivery efficiency. However, the optimization time reported in these studies [29, 28, 15]—ranging from several minutes to multiple hours—could pose a barrier to the clinical adoption of the framework. Based on the ground truth provided by modified ELO-SPAT, Ma et al. [15] demonstrates the feasibility of using supervised machine learning method to integrate features in geometric, WET and beamlet and generate ELs for each field. However, a supervised approach not only adds to the development workload due to ground truth generation but also constrains model performance to the quality of the pre-existing algorithms.
@@
A gap in the literature is the lack of deep learning-based approaches that can swiftly and directly predict optimal EL sequences, particularly using unsupervised learning. Deep learning has demonstrated strong potential in radiation oncology, offering high efficiency, automation, and generalization capacity [9]. Yet, its application to PAT planning, and especially to EL pre-selection, remains largely unexplored. Unsupervised learning offers a key advantage in this context: it eliminates the dependency on manually labeled data or algorithm-generated ground truths, enabling researchers to directly embed domain knowledge into customized loss functions that reflect clinical priorities such as dose coverage, OAR sparing, and delivery efficiency.
@@
In this study, we propose SPArcdl, an unsupervised deep learning model for fast and effective energy layer pre-selection in PAT plan optimization, which generates a data format specifically designed to support the application of deep learning models in EL pre-selection. These matrices serve as proxies for dose coverage and are analyzed by a U-Net-based neural network. The model is trained end-to-end using a tri-objective loss function that (1) maximizes target coverage, (2) minimizes OAR exposure, and (3) penalizes frequent and time-consuming EL switches.
@@
We evaluate SPArcdl on 54 NPC cases and compare its performance with SPArcps. Results show that SPArcdl significantly EL switch time—by 38.4%—while enhancing conformity index by $0.16$ ($p<0.01$), lowering homogeneity by $0.71$ ($p<0.01$), reducing the mean dose of brainstem by $0.21$ ($p<0.01$). By bypassing the need for supervised labels and offering rapid inference within $1$ second, SPArcdl demonstrates a viable path toward integrating deep learning into real-time PAT planning workflows, supporting broader clinical adoption through enhanced efficiency, automation, and plan quality.
@@ 2.1 Dataset and spot-count data representation
The treatment planning CT scans, acquired with a Philips Big Bore CT, of $54$ patients who has undergone nasopharyngeal carcinoma (NPC) radiotherapy were collected from our collaborative hospital [^1] Union Hospital, Tongji Medical College, Huazhong University of Science and Technology. Our method for pre-selecting ELs depends on the open-source software matRad (Department of Medical Physics in Radiation Oncology, German Cancer Research Center-DKFZ) [25] which uses the CTs to generate the initial complete range of ELs that longitudinally encompassed the entire treatment target at all gantry angles.
The inclusion of diverse information—such as beamlet, WET and geometric-related features [15]—may complicate the input structure of deep learning model and lead researchers to ambiguity. This study presents a novel data restructuring approach, termed spot-count representation, which constructs a data format specifically tailored to facilitate the application of deep learning models in EL pre-selection. The spot-count representation is founded on the assumption that the number of ELs, denoted by $E$, and the number of gantry angles, denoted by $G$, are finite. Accordingly, we construct a matrix $M\in\mathbb{R}^{G\times E}$, where rows correspond to sorted gantry angles and columns to sorted ELs, to record the number of proton spots intersecting a given anatomical structure. In this context, each element of $M$ represents the count of spots that intersect a specific structure $S$, organized by gantry angle and EL. For example, the number of spots intersecting CTV can be integrated into $M_{CTV}$.
Building on the concept of $M$, the number of spots intersecting one anatomical structure ($S_{1}$) but not another ($S_{2}$) can be represented by the element-wise subtraction $M_{S_{1}}-M_{S_{2}}$, resulting in a new matrix denoted as $M_{S_{1}-S_{2}}$. For instance, the number of spots that pass through the CTV while avoiding the brainstem is captured by the matrix $M_{CTV-Brainstem}$ ($M_{CTV-BRS}$). Since $M$ systematically records the number of spot intersections with anatomical structures, this data restructuring method is referred to as the spot-count representation.
Figure 1: (a) SPArcdl model architecture. (b) ELS time table displays the time cost when the gantry angle transitions from $g-1$ to $g$.
@@ 2.2 Method
We approximate the dose coverage of a volume as the number of spots intersection with the volume. Based on this approximation, matrix $M_{CTV}$ presents the dose coverage of CTV and $M_{CTV-BRS}$ presents dose covering CTV while dodging brainstem. Both matrices are formed and structured by gantry angles and EL. Previous method has found relative mean WET is key to mitigate the need for EL ascending and therefore improve delivery efficiency [17]. As shown in Figure 1, the mean value of WET for each gantry angle is calculated and fused with $M_{CTV}$ and $M_{CTV-BRS}$ by the gantry angle’s channel. These two matrices are then concatenated into a 2-channels tensor and fed to the network, $h$. In our experiments, the input is of size $2\times 72\times 103$, but the network is not limited to this specific size. This input size happens to be similar to the dimensionality of an image, which is channel $\times$ height $\times$ width. In the past decade, many deep learning architecture have been developed to solve computer vision tasks. We use $h$, a network similar to the classic image segmentation method U-Net  [18], to analyze the spot counts of CTV and brainstem in the context of ELs and gantry angles. The model consists of encoder and decoder sections with skip connection. We use convolution layers with a kernel size of $3$ and a stride of $1$ in both the encoder and decoder stages. Each convolutional layer is followed by a LeakyReLU activation function with a negative slope of $0.2$. The convolutional layers capture hierarchical spot-count distribution features of CTV and brainstem to estimate the optimal ELs. The encoder employs average pooling to reduce the spatial dimensions in half at each level. During the decoding stage, we alternate between upsampling, applying convolutions, and concatenating skip connections that transfer features learned in the encoding stages directly to the layers responsible for generating the EL selection. The spatial resolution of the encoder levels is one half of the previous level, and that of the decoder levels is twice of the previous level. Four levels in total are included in the network with depth of $16$, $32$, $64$, and $128$. We set $1$ as the channel number of network’s output and then perform Softmax on the logit vectors in gantry angle’s direction to obtain a probability tensor $P\in\mathbb{R}^{G\times EL}$. At inference stage, the EL corresponding to the highest probability in the logit vector for each gantry angle is selected as the final EL for that angle. We name this model, including the loss function, SPArcdl.
@@
The dataset is divided randomly into $5$ folds for cross-validation, and the reported results represent the mean validation performance across all folds. The proposed model is trained end-to-end, optimized with the AdamW [14] optimizer using a learning rate of $10^{-5}$ and a batch size of $24$. Training is conducted for $1000$ epochs. The model is implemented in PyTorch[16] and trained on a single NVIDIA GeForce 3090 GPU, and inference is run on a 11th Gen Intel(R) i7 CPU.
@@ 2.3 Unsupervised Loss Function
The three objectives of SPArcdl are to (1) maximize the dose coverage on CTV; (2) minimize the dose coverage on brainstem; and (3) reduce the energy shifting time across all gantry angles. The EL pre-selection task can be formulated as an EL classification problem at all gantry angles. As a popular loss function of classification problems, cross entropy provides a well-founded, interpretable, and computationally efficient way to measures the difference between two probability distributions. In this paper, the in-used cross entropy loss is formulated in Eq.(1).
@@
where $CE$ denotes cross entropy, $N$ denotes the number of patients, $EL$ denotes the number of energy layers, $G$ denotes the number of gantry angles, $M_{g,el}$ denotes the true labels for energy $el$ for angle $g$, and $P_{g,el}$ is the predicted probability for energy $el$ for angle $g$.
Specifically, to achieve objective (1), the first part of the loss function of Eq.(4), i.e. $CE(b(M_{CTV}),P)$, aims to shorten the distance between probability distribution of $P_{g}$ and $b(M_{CTV_{g}})$, so as to select the EL at angle $g$ that maximize the number of spots, or dose coverage, $g\in G$. $b$ is binarized-max function, which set the maximum value, or the maximum spot count, of each gantry angle as $1$ and other values as 00. Like-wise, to achieve objective (2), the second part of Eq.(4), i.e. $CE(b(M_{CTV-BRS}),P)$, is designed to maintain the coverage of CTV while reducing the brainstem dose exposure. The third part of Eq.(4) is introduced to penalize the ELS time, which can be generated using the ELS time table ($TB_{els}$) and formularized as Eq.(2).
@@
where $T(P_{g-1},P_{g},TB_{els})$ is the extraction of ELS time when the gantry angle switches from $g-1$ to $g$ given the time table $TB_{els}$. Figure 1 (b) illustrates $TB_{els}$ where both the vertical (top to bottom) and horizontal (left to right) axes represent EL indices sorted in ascending order based on EL values. Both dimensions have a length of 102, reflecting the total of $102$ distinct ELs presented in the dataset used. The cell value of $TB_{els}$ represents the ELS time between two adjacent gantry angles, $g_{-1}$ and $g$. It can be observed from Figure 1 (b) the time cost of ascended ELS is higher than that of descended ELS. Note the ELS time is zero when EL is unchanged. The total loss function is

### v2

@@ Abstract
Objective. Proton arc therapy (PAT) is an emerging and promising modality in radiotherapy, offering improved dose distribution and treatment robustness over conventional intensity-modulated proton therapy. Yet, identifying the optimal energy layer (EL) sequence remains challenging due to the intensive computational demand and prolonged in treatment delivery time. This study proposes an unsupervised deep learning framework for fast EL pre-selection that minimizes EL switch (ELS) time while maintaining high plan quality. Approach. We introduce a novel data representation method—spot-count representation—which encodes the number of proton spots intersecting the target and organs at risk (OAR) in a matrix structured by sorted gantry angles and energy layers. This representation serves as the input of an U-Net style architecture, SPArcdeep learning (SPArcdl), which is trained using a tri-objective function: maximizing spot-counts on target, minimizing spot-counts on OAR, and reducing ELS time. The model is evaluated on 35 nasopharyngeal cancer cases, and its performance is benchmarked against plans generated by SPArcparticle swarm. Main results. SPArcdl produces EL pre-selection that significantly improves both plan quality and delivery efficiency. Compared to SPArcparticle swarm, it enhances the conformity index by 0.1 ($p<0.01$), reduces the homogeneity index by 0.71 ($p<0.01$), lowers the brainstem mean dose by 0.25 ($p<0.01$), and shortens the ELS time by 37.2% ($p<0.01$). The results unintentionally reveal employing unchanged ELS is more time-wise efficient than descended ELS. SPArcdl’s inference time is within 1 second. However, SPArcdl plan demonstrates limitation in robustness. Significance. The proposed spot-count representation lays a foundation for incorporating unsupervised deep learning approaches into EL pre-selection task. SPArcdl is a fast tool for generating high-quality PAT plans by strategically pre-selecting EL to reduce delivery time while maintaining excellent dosimetric performance.
@@ 1 Introduction
A major approach to achieve the PAT planing optimization is to break it down into two steps: pre-selecting EL in advance, then using the selected EL to complete regular spot weight optimization [28]. In this approach, a significant challenge to the clinical adoption of PAT lies in the optimization of pre-selected energy layer (EL) sequences. In SPArc, the system is required to deliver protons at selected energy layers on each gantry angle. With hundreds of control points involved, frequent switching between energy layers can lead to substantial delays in treatment delivery. This is especially true for upward energy switches, which are time-intensive due to the magnetic hysteresis in the proton therapy system [8, 31, 2, 23, 22]. Consequently, minimizing energy layer switch (ELS) time without compromising dosimetric quality has become a focus in PAT plan optimization.
@@
Several EL sequence optimization approaches have been developed to address this challenge. Sanchez-Parcerisa et al. [19] proposes and compares four range optimization algorithms to select optimal mono- and bi-energetic beam energies for proton modulated arc therapy, aiming to minimize energy-switching time while maintaining dosimetric quality. However, this method shows constraints in performance when applied to intricate or irregular target geometries. The SPArc algorithm [6] improves dose delivery by iteratively refining gantry angles and redistributing EL. As a subsequent method, SPArcseq [12] reduces switching time by aligning energy sequences with the gantry’s rotational direction, and SPArcparticle swarm (SPArcps) [17] builds upon it by segmenting the arc into sectors based on water equivalent thickness (WET) and selecting EL that minimize switching time across control points. However, SPArcps may need more objective functions on organs at risk (OARs) to optimize Head and Neck cancer, including nasopharyngeal carcinoma (NPC). Additionally, the ELO-SPAT framework [29] integrates multiple objectives, including dose fidelity, energy sequencing constraints, and regularization terms, into a unified mathematical formulation that penalizes upward energy switches while exploiting downward transitions to enhance delivery efficiency. However, the optimization time reported in these studies [30, 29, 15]—ranging from several minutes to multiple hours—could pose a barrier to the clinical adoption of the framework. Based on the ground truth provided by modified ELO-SPAT, Ma et al. [15] demonstrates the feasibility of using supervised machine learning method to integrate features in geometric, WET and beamlet and generate EL for each field. However, a supervised approach not only adds to the development workload due to ground truth generation but also constrains model performance to the quality of the pre-existing algorithms.
@@
A gap in the literature is the lack of deep learning-based approaches, particularly using unsupervised learning, that can swiftly and directly predict optimal EL sequences. Deep learning has demonstrated strong potential in radiation oncology, offering high efficiency, automation, and generalization capacity [9]. Yet, its application to EL pre-selection for PAT planning optimization remains largely unexplored. Unsupervised learning offers key advantages in this context: firstly it eliminates the workload of manually labeled data or algorithm-generated ground truths; secondly it removes the performance constrain brought by the quality of the pre-existing algorithm that generates the ground truth; thirdly, it enables researchers to directly embed domain knowledge into customized loss functions that reflect clinical priorities such as dose coverage, OAR sparing, and delivery efficiency.
@@
In this study, we propose SPArcdeep learning (SPArcdl), an unsupervised deep learning model for fast and effective energy layer pre-selection in PAT plan optimization. SPArcdl takes in data generated by spot-count representation, a novel data representation method specifically designed to support the application of deep learning models in EL pre-selection. This input data serves as a proxy of dose coverage and are analyzed by a U-Net style neural network. The model produces an EL probability distribution at all gantry angles and it is optimized using a tri-objective loss function that (1) maximizes target coverage, (2) minimizes OAR dose exposure, and (3) penalizes frequent and time-consuming EL switch (ELS).
@@
We evaluate SPArcdl on 35 NPC cases and compare its performance with SPArcps. Results show that SPArcdl significantly reduces ELS time—by 37.2% (p<0.01p<0.01)—while enhancing conformity index (CI) by $0.1$ ($p<0.01$), lowering homogeneity index (HI) by $0.71$ ($p<0.01$), and reducing the mean dose of brainstem by $0.25$ ($p<0.01$). By bypassing the need for supervised labels and offering rapid inference within $1$ second, SPArcdl demonstrates a viable path toward integrating deep learning into PAT planning workflows, supporting broad clinical adoption through enhanced efficiency and plan quality.
@@ 2.1 Dataset and spot-count data representation
The treatment planning CT scans, acquired with a Philips Big Bore CT, of $35$ patients who has undergone NPC radiotherapy were collected from our collaborative hospital [^1] Union Hospital, Tongji Medical College, Huazhong University of Science and Technology. Our method for pre-selecting EL depends on the open-source software matRad 222developed by Department of Medical Physics in Radiation Oncology, German Cancer Research Center-DKFZ [26] which uses the CTs to generate the initial complete range of EL that longitudinally encompassed the entire treatment target at all gantry angles.
The inclusion of diverse information—such as beamlet, WET and geometric-related features [15]—may complicate the input structure of deep learning model. This study presents a novel data restructuring approach, termed spot-count representation, which constructs a data format specifically tailored to facilitate the application of deep learning models in EL pre-selection. The spot-count representation is founded on the assumption that the number of EL, denoted by $E$, and the number of gantry angles, denoted by $G$, are finite. Accordingly, we construct a matrix $M\in\mathbb{R}^{G\times E}$, where rows correspond to sorted gantry angles and columns to sorted EL, to record the number of proton spots intersecting a given anatomical structure. In this context, each element of $M$ represents spots count that intersect a specific structure $S$, organized by gantry angle and EL. For example, the number of spots intersecting CTV can be integrated into $M_{CTV}$.
Building on the concept of matrix $M$, the number of spots intersecting one anatomical structure ($S_{1}$) but not another ($S_{2}$) can be represented by the element-wise subtraction $M_{S_{1}}-M_{S_{2}}$, resulting in a new matrix denoted as $M_{S_{1}-S_{2}}$. For instance, the number of spots that pass through the CTV while avoiding the brainstem is captured by the matrix $M_{CTV-Brainstem}$ ($M_{CTV-BRS}$). Since $M$ systematically records the number of spot intersections with anatomical structures, this data restructuring method is referred to as the spot-count representation.
Figure 1: (a) SPArcdl model architecture. (b) ELS time table displays the time cost when the gantry angle transitions from g1g-1 to $g$.
@@ 2.2 Method
We approximate the dose coverage of a volume as the number of spots intersection with the volume. Based on this approximation, matrix $M_{CTV}$ presents the dose coverage of CTV and $M_{CTV-BRS}$ presents dose covering CTV while dodging brainstem. Both matrices are formed and structured by gantry angles and EL. Previous method has found the WET value is key to mitigate the need for EL ascending and therefore improve delivery efficiency [17]. As shown in Figure 1, the max value of WET of each gantry angle is calculated and fused with $M_{CTV}$ and $M_{CTV-BRS}$ by the gantry angle’s channel. These two matrices are then concatenated into a 2-channels tensor and fed to the network, $h$. In our experiments, the input is of size $2\times 72\times 103$, but the network is not limited to this specific size. This input dimension happens to be similar to the dimensionality of an image, which is channel $\times$ height $\times$ width. In the past decade, many deep learning architecture have been developed to solve computer vision tasks. We use $h$, a network similar to the classic image segmentation method U-Net  [18], to analyze the spot counts of CTV and brainstem in the context of EL and gantry angles. The model consists of encoder and decoder sections with skip connection. We use convolution layers with a kernel size of $3$ and a stride of $1$ in both the encoder and decoder stages. Each convolutional layer is followed by a LeakyReLU activation function with a negative slope of $0.2$. The convolutional layers capture hierarchical spot-count distribution features of CTV and brainstem to estimate the optimal EL. The encoder employs average pooling to reduce the spatial dimensions in half at each level. During the decoding stage, we alternate between upsampling, applying convolutions, and concatenating skip connections that transfer features learned in the encoding stages directly to the layers responsible for generating the EL selection. The spatial resolution of the encoder levels is one half of the previous level, and that of the decoder levels is twice of the previous level. Four levels in total are included in the network with depth of $16$, $32$, $64$, and $128$. We set $1$ as the channel number of network’s output which is of size $2\times 72\times 102$, and perform Softmax along the gantry angle’s direction to obtain a probability tensor $P\in\mathbb{R}^{G\times EL}$. At inference stage, the EL corresponding to the highest probability of each gantry angle is selected as the predicted EL for that angle. We name this model, including the loss function, SPArcdl.
@@
The dataset is divided randomly into $5$ folds for cross-validation, and the reported results represent the mean validation performance across all folds. The proposed model is trained end-to-end, optimized with the AdamW optimizer[14] using a learning rate of $10^{-5}$ and a batch size of $16$. Training is conducted for $1000$ epochs. The model is implemented in PyTorch[16] and trained on a single NVIDIA GeForce 3090 GPU, and inference is run on 11th Gen Intel(R) i7 CPU.
@@ 2.3 Unsupervised loss function
Spot count is used as a proxy for dose coverage in SPArcdl. The three objectives of SPArcdl are to (1) maximize the dose coverage, or spot count, on CTV; (2) minimize the dose coverage, or spot count, on brainstem; and (3) reduce the ELS time across all gantry angles. The EL pre-selection task can be formulated as an EL classification problem at all angles. As a popular loss function of classification problems, cross entropy provides a well-founded, interpretable, and computationally efficient way to measures the difference between two probability distributions. In this paper, the in-used cross entropy loss is formulated in Eq.(1).
@@
where $CE$ denotes cross entropy, $N$ denotes the case number of a batch, $EL$ denotes the number of energy layers, $G$ denotes the number of gantry angles, $M_{g,el}$ denotes the true labels for energy $el$ for angle $g$, and $P_{g,el}$ is the predicted probability for energy $el$ for angle $g$.
Specifically, to achieve objective (1), the first part of the loss function of Eq.(4), i.e. $CE(b(M_{CTV}),P)$, aims to shorten the distance between probability distribution of $P_{g}$ and $b(M_{CTV_{g}})$, so as to select the EL at angle $g$ that maximize the number of spots, or dose coverage, $g\in G$. $b$ is binarized-max function, which set the maximum value, or the maximum spot count, of each gantry angle as $1$ and other values as 0. Like-wise, to achieve objective (2), the second part of Eq.(4), i.e. $CE(b(M_{CTV-BRS}),P)$, is designed to maintain the coverage of CTV while reducing the brainstem dose exposure. The third part of Eq.(4) is introduced to penalize the ELS time, which can be generated using the ELS time table ($TB_{els}$) and formularized as Eq.(2).
@@
where $T(P_{g-1},P_{g},TB_{els})$ is the extraction of ELS time when the gantry angle switches from g1g-1 to $g$ given the time table $TB_{els}$. Figure 1 (b) illustrates $TB_{els}$ where both the vertical (top to bottom) and horizontal (left to right) axes represent EL indices sorted in ascending order based on EL values. Both dimensions have a length of 102, reflecting the total of $102$ distinct EL presented in the dataset used. The cell value of $TB_{els}$ represents the ELS time between two adjacent gantry angles, $g_{-1}$ and $g$. It can be observed from Figure 1 (b) the time cost of ascended ELS is higher than that of descended ELS. Note the ELS time is zero when EL is unchanged. The total loss function is